1. A tennis ball traveling at 12 m/s was approaching the tennis player’s racket. The ball was hit back with a speed of 38 m/s in the opposite direction from where it came.


a. Compute the change in the momentum of the ball.

b. Suppose the mass of the tennis ball is 57 g. If it came in contact with the racket for 0.03 second, what is the average force exerted on the ball?

1
Expert's answer
2021-04-13T06:40:22-0400

a. Compute the change in the momentum of the ball:


"\Delta p=m(v_f-v_i)."

b. The average force would be


"F=\frac{\Delta p}{t}=\frac{m(v_f-v_i)}{t},\\\space\\
F=\frac{0.057(38-(-12))}{0.03}=95\text{ N}."